# Reachability for multiagent control systems via Lyapunov functions

### v1

@@ 1. Introduction
$\dot{x}(t)\in F(x(t),\mu_{t})\mbox{ for almost every $t\geq 0$, }$
where $F:\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\rightrightarrows\mathbb{R}^{d}$ is a set-valued map and $\mathcal{P}_{2}(\mathbb{R}^{d})$ is the set of probability measures on $\mathbb{R}^{d}$ with finite second moment. Namely, the velocity of each agent depends both on his own position and on the position of the whole crowd of agents. It is worth pointing out that we make an indistiguishability assumption meaning that an agent does not interact individually with every other agents but only with the whole crowd of agents.
@@
where $v_{t}(\cdot)$ is a Borel time dependent vector field on $\mathbb{R}^{d}$. This equation holds true in the sense of distributions on ×$\mathbb{R}\times\mathbb{R}^{d}$ (cf. Definition 2.2 later on).
@@
The main results of the paper are contained in Theorems 4.10 and 4.16. We investigate two rates of convergence of the trajectory: either asymptotically with an exponential decay or linearly in finite time horizon. A key ingredient in our approach is a new comparison result for Hamilton Jacobi equations in Wasserstein spaces (a slightly different version of the result developed in [JMQ20]), which is a byproduct of our article.
The paper is organized as follows: a preliminary section is devoted to notations, essential recalls on tools of optimal transport theory and introduction of the dynamical setting under investigation. In Section 3, we recall basic facts on Hamilton Jacobi Equations and we state an adapted version of the comparison principle for viscosity solutions on the Wasserstein space, whose proof is postponed to Appendix B. Section 4 contains the core of the paper. Finally, Appendix A is devoted to some technical proofs.
@@ 2.1. The Wasserstein metric
Let $X$ be a separable Banach space (here $X$ will be either $\mathbb{R}^{d}$ or $\mathrm{C}([a,b];\mathbb{R}^{d})$, the space of $\mathbb{R}^{d}$-valued continuous curves on $[a,b]\subset\mathbb{R}$ endowed with the uniform norm). We denote by $\mathcal{P}(X)$ the space of Borel probability measures on $X$, endowed with the topology of narrow convergence111 We recall that a sequence $\{\nu_{n}\}_{n\in\mathbb{N}}\subset\mathcal{P}(X)$ narrowly converges to $\nu\in\mathcal{P}(X)$, and we write $\nu_{n}\rightharpoonup\nu$ as $n\to+\infty$, if for any $f\in\mathrm{C}([a,b];\mathbb{R})$ bounded $\int_{X}f(x)\,\mathrm{d}\nu_{n}(x)\to\int_{X}f(x)\,\mathrm{d}\nu(x),\quad\text{as }n\to+\infty.$
@@ 2.2. Trajectories of the control system
for any $(x,\nu)\in\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})$, where $C:=\max\{1,L\,K_{\mathrm{F}}\}$.
@@ 3. Hamilton Jacobi Bellman equation
We introduce the following Hamilton-Jacobi-Bellman equation in $[a,b]\times\mathcal{P}_{2}(\mathbb{R}^{d})$ (with $a<b$),
@@ 3.1. Sub/super-differentials
Definition 3.3 (Sub/super-differential in ×$\mathbb{R}\times\mathcal{P}_{2}(\mathbb{R}^{d})$).
Let $a,b\in\mathbb{R}$, $a<b$, and consider the map $U:[a,b]\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$. Given $(\bar{t},\bar{\nu})\in[a,b]\times\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\delta\geq 0$, we say that the pair $(p_{\bar{t}},p_{\bar{\nu}})\in\mathbb{R}\times L^{2}_{\bar{\nu}}(\mathbb{R}^{d})$ is a (total) $\delta$-viscosity subdifferential of $U$ at $(\bar{t},\bar{\nu})$, and we write $(p_{\bar{t}},p_{\bar{\nu}})\in\boldsymbol{D}^{-}_{\delta}U(\bar{t},\bar{\nu})$, if
@@
    $U(t,\nu)-U(\bar{t},\bar{\nu})\geq p_{\bar{t}}\,(t-\bar{t})+\int_{\mathbb{R}^{d}}\langle p_{\bar{\nu}}(x),y-x\rangle\,\mathrm{d}\boldsymbol{\sigma}(x,y)-\delta\cdot\Delta_{t,\boldsymbol{\sigma}}+o\left(\Delta_{t,\boldsymbol{\sigma}}\right),$
@@
    $U(\nu)-U(\bar{\nu})\geq\int_{\mathbb{R}^{d}}\langle p(x),y-x\rangle\,\mathrm{d}\boldsymbol{\sigma}(x,y)-\delta\cdot\Delta_{\boldsymbol{\sigma}}+o\left(\Delta_{\boldsymbol{\sigma}}\right),$
@@
Let $a,b\in\mathbb{R}$, $a<b$, $(\bar{t},\bar{\nu})\in(a,b)\times\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\delta\geq 0$. Let $U:[a,b]\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be partially differentiable w.r.t. $t\in(a,b)$ at $(\bar{t},\bar{\nu})$, so that there exists $\partial_{t}U(\bar{t},\bar{\nu})$. Then
@@
(3.5) $U(t,\nu)-U(\bar{t},\bar{\nu})\geq p_{\bar{t}}\,(t-\bar{t})+\int_{\mathbb{R}^{d}}\langle p_{\bar{\nu}}(x),y-x\rangle\,\mathrm{d}\boldsymbol{\sigma}(x,y)-\delta\cdot\Delta_{t,\boldsymbol{\sigma}}+o\left(\Delta_{t,\boldsymbol{\sigma}}\right).$
@@ 3.2. Viscosity solutions and the Comparison principle
We state here the Comparison Principle for (3.1) that will be used in this paper. Its proof is postponed to Appendix B.
Theorem 3.8 (Comparison).
  1. (1)
  2. (2)
    $|u_{i}(t,\nu_{1})-u_{i}(t,\nu_{2})|\leq L_{R,i}\,W_{2}(\nu_{1},\nu_{2}),$
  3. (3)
    (3.6) $\partial_{t}U(t,\nu)+\mathcal{H}(\nu,D_{\nu}U(t,\nu))=0$
4. Reachability through Lyapunov functions
4.1. Lyapunov function and asymptotic reachability
We introduce Lyapunov functions for a suitable Lipschitz continuity notion that we call $\mathsf{m}_{2}$-local Lipschitzianity.
@@ 4.1. Lyapunov function and asymptotic reachability
To carry on our main study of Section 4.2, we are interested in the following function $\varphi:[0,+\infty)\to\mathbb{R}$, that will be associated with the characterization of asymptotic reachability (see next Definition 4.5):
@@
where $\alpha>0$ is a given parameter. Analogously, in Section 4.3 we will present the case
(4.4) $\varphi(y):=\alpha,$
with $\alpha>0$ given; this one will be involved, instead, in the characterization of reachability.
@@
  1. (1)
    $\nu$ is reachable if for any $\mu_{0}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ there exist $T>0$ and $\mu\in\mathcal{A}_{[0,T]}(\mu_{0})$ such that $\mu_{t=T}=\nu$.
@@ 4.2. Main results
In the whole section, we consider $\varphi$ as in (4.3) with $\alpha>0$ a given parameter.
@@
Let $\mathrm{V}:\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,+\infty)$ be a $\mathsf{m}_{2}$-locally Lipschitz map and $T>0$. The map $\mathrm{V}$ is a Lyapunov function for $\mathrm{F}$ with rate $\varphi$ if and only if the function $\mathrm{S}:[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,+\infty)$, defined by $\mathrm{S}(t,\nu):=e^{\alpha\,t}\mathrm{V}(\nu)$, is a partial-viscosity supersolution of
@@
Let $T>0$ and $\mathrm{V}:\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,+\infty)$ be a Lyapunov function for $\mathrm{F}$ with rate $\varphi$, namely $\mathrm{V}$ is $\mathsf{m}_{2}$-locally Lipschitz and there exists a map $C:\mathcal{P}_{2}(\mathbb{R}^{d})\to(0,+\infty)$ bounded on bounded sets such that the following holds
@@
By the arbitrariness of $T>0$ and since the gluing of admissible trajectories is still an admissible trajectory, we conclude that there exists $\mu\in\mathcal{A}_{[0,+\infty)}(\mu_{0})$ such that the map $SS\SS$, is nonincreasing in $[0,+\infty)$. This proves the left-to-right implication in Proposition 4.8.
@@
00 0 $\geq\mathrm{S}(t,\mu_{t})-\mathrm{S}(\bar{t},\bar{\nu})=e^{\alpha t}\mathrm{V}(\mu_{t})-e^{\alpha\bar{t}}\mathrm{V}(\bar{\nu})$
@@
(4.14) $\liminf_{t\to\bar{t}^{+}}\frac{\Delta_{\boldsymbol{\pi}_{t}}}{t-\bar{t}}\left(\delta-\frac{o(\Delta_{\boldsymbol{\pi}_{t}})}{\Delta_{\boldsymbol{\pi}_{t}}}\right)\geq\alpha V(\bar{\nu})+\liminf_{t\to\bar{t}^{+}}\frac{1}{t-\bar{t}}\int\langle p_{\bar{\nu}}(x),y-x\rangle\,\mathrm{d}\boldsymbol{\pi}_{t}(x,y).$
@@
Following [Au, AuF], we can give another characterization for Lyapunov functions in terms of viability conditions over their epigraph, as follows.
@@
Assume Hypothesis 2.1. Let $\mathrm{V}:\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,+\infty)$ be a $\mathsf{m}_{2}$-locally Lipschitz map. Then $\mathrm{V}$ is a Lyapunov function for $\mathrm{F}$ with rate $\varphi$ if and only if $\mathrm{epi}\,(\mathrm{V})$ is viable for the coupled dynamics
@@
[…]
@@
[…]
@@
Assume Hypothesis 2.1. Let $\bar{\nu}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and assume there exists a Lyapunov function $\mathrm{V}$ for $\mathrm{F}$ with rate $\varphi$ such that $\mathrm{V}(\bar{\nu})=0$, $\mathrm{V}(\nu)>0$ for any $\nu\neq\bar{\nu}$ and satisfying either one of the following conditions:
@@
    Assume (3) holds. Then, (4.20) implies that there exists $\tilde{K}>0$ such that $\mathsf{m}_{2+\varepsilon}(\mu_{t})\leq\tilde{K}$ for all $t\geq 0$. Reasoning as in item (1), there exists $\{t_{n}\}_{n\in\mathbb{N}}$ and $\bar{\mu}\in\mathcal{P}(\mathbb{R}^{d})$ such that $t_{n}\to+\infty$ as $n\to+\infty$ and $\mu_{t_{n}}\rightharpoonup\bar{\mu}$ narrowly as $n\to+\infty$. Moreover, $\{\mu_{t}\}_{t\geq 0}$ has uniformly integrable $2$-moments (see [ags, Equation (5.1.20)]), hence $\mu_{t_{n}}\overset{W_{2}}{\to}\bar{\mu}$ as $n\to+\infty$. We conclude by the continuity of $V$ in $W_{2}$-metric.
4.3. The case $\varphi(y)=\alpha$
In this section, similarly as for Section 4.2, we study sufficient conditions for (finite-time) reachability via Lyapunov functions. Differently from Section 4.2, throughout this section we consider $\varphi$ as in (4.4), i.e. $\varphi(y):=\alpha,$ with $\alpha>0$ a given parameter.
Unless otherwise explicitly stated, the proofs follow the same track as the corresponding results in Section 4.2, therefore they are omitted.
Lemma 4.11.
Let $\mathrm{V}:\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,+\infty)$ be a $\mathsf{m}_{2}$-locally Lipschitz map and $T>0$. The map $\mathrm{V}$ is a Lyapunov function for $\mathrm{F}$ with rate $\varphi$ if and only if the function $\mathrm{S}:[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,+\infty)$, defined by $\mathrm{S}(t,\nu):=\alpha\,t+\mathrm{V}(\nu)$, is a partial-viscosity supersolution of
Proof.
It is sufficient to notice that $\partial_{\delta}^{-}\mathrm{S}(t,\nu)=\partial_{\delta}^{-}\mathrm{V}(\nu)$ for any $(t,\nu)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})$. ∎
Lemma 4.12.
Let $T>0$ and $\mathrm{V}:\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,+\infty)$ be a Lyapunov function for $\mathrm{F}$ with rate $\varphi$, namely $\mathrm{V}$ is $\mathsf{m}_{2}$-locally Lipschitz and there exists a map $C:\mathcal{P}_{2}(\mathbb{R}^{d})\to(0,+\infty)$ bounded on bounded sets such that the following holds
(4.23) $\alpha+\mathcal{H}_{\mathrm{F}}(\nu,p)\leq C(\nu)\delta,\qquad\forall\,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d}),\,\forall\,p\in\partial_{\delta}^{-}\mathrm{V}(\nu),\,\forall\,\delta>0.$
Then the function $\mathrm{S}:[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,+\infty)$, defined by $\mathrm{S}(t,\nu):=\alpha t+\mathrm{V}(\nu)$, is a viscosity supersolution of
in the sense of Definition 3.7.
Proof.
Proceding as in the proof of Lemma 4.7, we easily get that
$p_{t}+\mathcal{H}_{\mathrm{F}}(\nu,p_{\nu})\leq(C(\nu)+1)\delta,$
for any $(t,\nu)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})$, $(p_{t},p_{\nu})\in\boldsymbol{D}^{-}_{\delta}\mathrm{S}(t,\nu)$, $\delta>0$, which gives the thesis. ∎
Proposition 4.13.

### v2

@@ 1. Introduction
$\dot{x}(t)\in\mathrm{F}(x(t),\mu_{t})\mbox{ for almost every $t\geq 0$, }$
where $\mathrm{F}:\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\rightrightarrows\mathbb{R}^{d}$ is a set-valued map and $\mathcal{P}_{2}(\mathbb{R}^{d})$ is the set of probability measures on $\mathbb{R}^{d}$ with finite second moment. Namely, the velocity of each agent depends both on his own position and on the position of the whole crowd of agents. It is worth pointing out that we make an indistiguishability assumption meaning that an agent does not interact individually with every other agents but only with the whole crowd of agents.
@@
where $v_{t}(\cdot)$ is a Borel time dependent vector field on $\mathbb{R}^{d}$. This equation holds true in the sense of distributions on $\mathbb{R}\times\mathbb{R}^{d}$ (cf. Definition 2.2 later on).
@@
The main results of the paper are contained in Theorem 4.7. In particular, we investigate an asymptotic rate of convergence of the trajectory with an exponential decay. A key ingredient in our approach is a comparison result for Hamilton Jacobi equations in Wasserstein spaces, which is a byproduct of our article. Unlike the analogous result developed in [JMQ20], which serves as inspiration for our approach, we do not assume boundedness of the viscosity solutions involved. Instead, we show that the result can be established under a suitable Lipschitz continuity condition, which fits more naturally within our framework.
The paper is organized as follows: a preliminary section is devoted to notations, essential recalls on tools of optimal transport theory and introduction of the dynamical setting under investigation. In Section 3, we recall basic facts on Hamilton Jacobi Equations and we state an adapted version of the comparison principle for viscosity solutions on the Wasserstein space, whose proof is postponed to Appendix B. Section 4 contains the core of the paper; examples of application of the theory are presented in Section 5. Finally, Appendix A is devoted to some technical proofs.
@@ 2.1. The Wasserstein metric
Let $X$ be a separable Banach space (here $X$ will be either $\mathbb{R}^{d}$ or $\mathrm{C}([a,b];\mathbb{R}^{d})$, the space of $\mathbb{R}^{d}$-valued continuous curves on $a,b]\subset\mathbb{R}$ endowed with the uniform norm). We denote by $\mathcal{P}(X)$ the space of Borel probability measures on $X$, endowed with the topology of narrow convergence [^1] We recall that a sequence $\{\nu_{n}\}_{n\in\mathbb{N}}\subset\mathcal{P}(X)$ narrowly converges to $\nu\in\mathcal{P}(X)$, and we write $\nu_{n}\rightharpoonup\nu$ as $n\to+\infty$, if for any $f\in\mathrm{C}([a,b];\mathbb{R})$ bounded $\int_{X}f(x)\,\mathrm{d}\nu_{n}(x)\to\int_{X}f(x)\,\mathrm{d}\nu(x),\quad\text{as }n\to+\infty.$ .
@@ 2.2. Trajectories of the control system
for any $(x,\nu)\in\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})$, where $C:=\max\{L,K_{\mathrm{F}}\}$.
@@ 3. Hamilton Jacobi Bellman equation
We introduce the following Hamilton-Jacobi-Bellman equation in $a,b]\times\mathcal{P}_{2}(\mathbb{R}^{d})$ (with $a<b$),
@@ 3.1. Sub/super-differentials
Definition 3.3 (Sub/super-differential in $\mathbb{R}\times\mathcal{P}_{2}(\mathbb{R}^{d})$).
Let $\in\mathbb{R}, a<ba<b$, and consider the map $U:[a,b]\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$. Given $(\bar{t},\bar{\nu})\in[a,b]\times\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\delta\geq 0$, we say that the pair $(p_{\bar{t}},p_{\bar{\nu}})\in\mathbb{R}\times L^{2}_{\bar{\nu}}(\mathbb{R}^{d})$ is a (total) $\delta$-viscosity subdifferential of $U$ at $(\bar{t},\bar{\nu})$, and we write $(p_{\bar{t}},p_{\bar{\nu}})\in\boldsymbol{D}^{-}_{\delta}U(\bar{t},\bar{\nu})$, if
@@
    $U(t,\nu)-U(\bar{t},\bar{\nu})\geq p_{\bar{t}}\,(t-\bar{t})+\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\langle p_{\bar{\nu}}(x),y-x\rangle\,\mathrm{d}\boldsymbol{\sigma}(x,y)-\delta\cdot\Delta_{t,\boldsymbol{\sigma}}+o\left(\Delta_{t,\boldsymbol{\sigma}}\right),$
@@
    $U(\nu)-U(\bar{\nu})\geq\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\langle p(x),y-x\rangle\,\mathrm{d}\boldsymbol{\sigma}(x,y)-\delta\cdot\Delta_{\boldsymbol{\sigma}}+o\left(\Delta_{\boldsymbol{\sigma}}\right),$
@@
Let $\in\mathbb{R}, a<ba<b$, $(\bar{t},\bar{\nu})\in(a,b)\times\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\delta\geq 0$. Let $U:[a,b]\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be partially differentiable w.r.t. $t\in(a,b)$ at $(\bar{t},\bar{\nu})$, so that there exists $\partial_{t}U(\bar{t},\bar{\nu})$. Then
@@
(3.5) $U(t,\nu)-U(\bar{t},\bar{\nu})\geq p_{\bar{t}}\,(t-\bar{t})+\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\langle p_{\bar{\nu}}(x),y-x\rangle\,\mathrm{d}\boldsymbol{\sigma}(x,y)-\delta\cdot\Delta_{t,\boldsymbol{\sigma}}+o\left(\Delta_{t,\boldsymbol{\sigma}}\right).$
@@ 3.2. Viscosity solutions and the Comparison principle
We state here in Theorem 3.11 the Comparison Principle for (3.1) that will be used in this paper and whose proof is postponed to Appendix B. The result involves a suitable Lipschitz continuity notion that we call $\mathsf{m}_{2}$-local Lipschitzianity.
@@
Theorem 3.11 (Comparison).
  1. (1)
  2. (2)
    $|u_{i}(t,\nu_{1})-u_{i}(t,\nu_{2})|\leq L_{R,i}\,W_{2}(\nu_{1},\nu_{2}),$
  3. (3)
    (3.7) $\partial_{t}U(t,\nu)+\mathcal{H}(\nu,D_{\nu}U(t,\nu))=0$
4. Reachability through Lyapunov functions
4.1. Lyapunov function and asymptotic reachability
We introduce Lyapunov functions satisfying the $\mathsf{m}_{2}$-local Lipschitzianity property introduced in Definition 3.8.
@@ 4.1. Lyapunov function and asymptotic reachability
In the whole paper, we consider the following function $\varphi:[0,+\infty)\to\mathbb{R}$, that will be associated with the characterization of asymptotic reachability (see next Definition 4.2):
@@
where $\alpha>0$ is a given parameter.
@@
Let $\mathrm{V}:\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,+\infty)$ be a $\mathsf{m}_{2}$-locally Lipschitz map and $T>0$. The map $\mathrm{V}$ is a Lyapunov function for $\mathrm{F}$ with rate $\varphi$ as in (4.2) if and only if the function $\mathrm{S}:[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,+\infty)$, defined by $\mathrm{S}(t,\nu):=e^{\alpha\,t}\mathrm{V}(\nu)$, is a partial-viscosity supersolution of
@@
Let $T>0$ and $\mathrm{V}:\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,+\infty)$ be a Lyapunov function for $\mathrm{F}$ with rate $\varphi$ as in (4.2), namely $\mathrm{V}$ is $\mathsf{m}_{2}$-locally Lipschitz and there exists a map $C:\mathcal{P}_{2}(\mathbb{R}^{d})\to(0,+\infty)$ bounded on bounded sets such that the following holds
@@
By the arbitrariness of $T>0$ and since the gluing of admissible trajectories is still an admissible trajectory, we conclude that there exists $\mu\in\mathcal{A}_{[0,+\infty)}(\mu_{0})$ such that the map $S\SS$, is nonincreasing in $0,+\infty)$. This proves the left-to-right implication in Proposition 4.5.
@@
$0$ $\geq\mathrm{S}(t,\mu_{t})-\mathrm{S}(\bar{t},\bar{\nu})=e^{\alpha t}\mathrm{V}(\mu_{t})-e^{\alpha\bar{t}}\mathrm{V}(\bar{\nu})$
@@
(4.12) ${\color[rgb]{0,0,0}\limsup_{t\to\bar{t}^{+}}}\frac{\Delta_{\boldsymbol{\pi}_{t}}}{t-\bar{t}}\left(\delta-\frac{o(\Delta_{\boldsymbol{\pi}_{t}})}{\Delta_{\boldsymbol{\pi}_{t}}}\right)\geq\alpha V(\bar{\nu})+\liminf_{t\to\bar{t}^{+}}\frac{1}{t-\bar{t}}\int\langle p_{\bar{\nu}}(x),y-x\rangle\,\mathrm{d}\boldsymbol{\pi}_{t}(x,y).$
@@
Following [Au, AuC, AuF], we can give another characterization for Lyapunov functions in terms of viability conditions over their epigraph, as follows.
@@
Assume Hypothesis 2.1. Let $\mathrm{V}:\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,+\infty)$ be a $\mathsf{m}_{2}$-locally Lipschitz map. Then $\mathrm{V}$ is a Lyapunov function for $\mathrm{F}$ with rate $\varphi$ as in (4.2) if and only if $\mathrm{epi}\,(\mathrm{V})$ is viable for the coupled dynamics
@@
Notice that condition $(\mathrm{V1})$ is equivalent to the following one, where the initial condition is taken only at the boundary of $\mathrm{epi}\,(\mathrm{V})$:
@@
Thanks to Proposition 4.5, to conclude we only need to prove that $(\mathrm{V2})$ is equivalent to
@@
Clearly (4.16) implies $(\mathrm{V2})$. Conversely, let us assume that $(\mathrm{V2})$ holds true. Fix $\mu_{0}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $T>0$. Consider for any integer $n>0$ the subdivision of $[0,T]$ given by $k\frac{T}{n}$, for $k=0,1,\ldots,n$. By $(\mathrm{V2})$ there exists $\mu^{n}\in\mathcal{A}_{[0,\frac{T}{n}]}(\mu_{0})$ such that
@@
Applying again $(\mathrm{V2})$ with the initial condition $\mu^{n}_{\frac{T}{n}}$, we can extend $\mu^{n}$ on $\left[\frac{T}{n},2\frac{T}{n}\right]$ such that
@@
Assume Hypothesis 2.1. Let $\bar{\nu}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and assume there exists a Lyapunov function $\mathrm{V}$ for $\mathrm{F}$ with rate $\varphi$ as in (4.2) such that $\mathrm{V}(\bar{\nu})=0$, $\mathrm{V}(\nu)>0$ for any $\nu\neq\bar{\nu}$ and satisfying either one of the following conditions:
@@
    Assume (3) holds. Then, (4.18) implies that there exists $\tilde{K}>0$ such that $\mathsf{m}_{2+\varepsilon}(\mu_{t})\leq\tilde{K}$ for all $t\geq 0$. Reasoning as in item (1), there exists $\{t_{n}\}_{n\in\mathbb{N}}$ and $\bar{\mu}\in\mathcal{P}(\mathbb{R}^{d})$ such that $t_{n}\to+\infty$ as $n\to+\infty$ and $\mu_{t_{n}}\rightharpoonup\bar{\mu}$ narrowly as $n\to+\infty$. Moreover, $\{\mu_{t}\}_{t\geq 0}$ has uniformly integrable 22-moments (see [ags, Equation (5.1.20)]), hence $\mu_{t_{n}}\overset{W_{2}}{\to}\bar{\mu}$ as $n\to+\infty$. We conclude by the continuity of $V$ in $W_{2}$-metric.
5. Applications
In this section we give two examples of application of our main result.
5.1. The second moment as a Lyapunov function
Let $\alpha>0$ and consider the following set-valued vector field $\mathrm{F}:\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\rightrightarrows\mathbb{R}^{d}$ driving the dynamics in Definition 2.2,
$\mathrm{F}(x,\nu):=\overline{B\big(0,\alpha\left(|x|+\mathsf{m}_{2}(\nu)\right)\big)},$
so that $\mathrm{F}$ satisfies Hypothesis 2.1. We can prove that $\bar{\nu}:=\delta_{0}$ is weakly asymptotically reachable thanks to Theorem 4.7. Indeed, we can show that the function $\mathrm{V}:\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,+\infty)$, defined by $\mathrm{V}(\nu):=\frac{1}{2}\mathsf{m}_{2}^{2}(\nu)$, is a Lyapunov function for $\mathrm{F}$ with rate $\varphi$ as in (4.2), satisfying the required conditions. In particular, by Proposition 4.5, in order to prove that $\mathrm{V}$ is a Lyapunov function for $\mathrm{F}$ with rate $\varphi$ it is sufficient to prove the following claims:
  • Claim 1: $\mathrm{V}$ is $\mathsf{m}_{2}$-locally Lispchitz.